Suppose​ 36% of​ movie-goers say that horror is their favorite movie genre. If a random

sample of 120​ movie-goers is​ taken, what are the mean and standard deviation of the

number of​ movie-goers that prefer​ horror?

A.

mu equals 0.36 comma sigma equals 0.0438μ = 0.36, σ = 0.0438

B.

mu equals 0.36 comma sigma equals 0.0019μ = 0.36, σ = 0.0019

C.

mu equals 43.2 comma sigma equals 27.648μ = 43.2, σ = 27.648

D.

mu equals 43.2 comma sigma equals 5.2581
Answer #1

Solution :

Given that,

p = 36%=0.36

q = 1 - p =1-0.36=0.64

n = 120

Using binomial distribution,

= n * p = 120*0.36=43.2

= n * p * q = 120*0.36*0.64=5.2581

D.

mu equals 43.2 comma sigma equals 5.2581